Find the molarity of 45.0 g NaC1 in 500.0 mL


1
Expert's answer
2022-05-23T17:01:03-0400

CM = n / V

C– concentration of molarity (mol/l);

n – amount substance of solute (mol);

V – volume of solution (L).

1. We find the amount of substance (n) of 45 gr NaCl. To do this, we divide the mass of NaCl (m = 45 g) by its molecular mass (MNaCl =C 58,5 g / mol):

n= m/M = 45 / 58,5 = 0,77 mol.

2. CM = n/V= 0,77 mol / 0,5litres = 1,54mol/l

ANSWER: 1,54 mol/l